Polar coordinates: 13. Describe the following curves: a./ r = 3, b./ r cos 0 = 3 c./ r sin 0 = 3 14. Convert from polar to reIf you can show work to prove your question that would be greatly appreciated.
Answer #1

Party We know 2 = 91 Coso , y=91 Sino, a) 9=2 Sino 9(91) = 2 9 Sino gr = 24 (artys_2y = 0 b) 9= 2 COSO g1Y = 299 caso rtyr =c) 9= Gos - Sino 810090 - 9 sino = 1 2-4 = 1 | x-y-1=0. Thank you